OpenStudy (anonymous):

Identify the values of a, b, and c in the following quadratic equation. 2x^2 + 3x + 7 = 0

Parth (parthkohli):

\[ax^2 + bx + c = 0\]\[2x^2 + 3x + 7 = 0 \]Hinty hinty hint.

OpenStudy (anonymous):

a=7 b=2 c=3?

Parth (parthkohli):

Hmm. You did the opposite thing.

OpenStudy (eyust707):

Take a look at the two equations... \(2x^2 + 3x + 7 = 0 \) \(ax^2 + bx + c = 0\) a, b and c refer to the coefficients. What number is in the place of a? What number is in the place of b? What number is in the place of c?

OpenStudy (anonymous):

a=3,b=2,c=7...

OpenStudy (eyust707):

Now you are just guessing

Parth (parthkohli):

Nono. I believe that a = 2.

OpenStudy (eyust707):

The 2 is in the a's spot!!

Parth (parthkohli):

\[\color{green}{a}x^2 + \color{red}{b}x + \color{yellow}c = 0 \\ \color{green}{2}x^2 + \color{red}{3}x + \color{yellow}{7}=0\]

OpenStudy (anonymous):

hmm.. a=2,b=3,c=7?

OpenStudy (anonymous):

ohhok.

Parth (parthkohli):

Now that's correct.